OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x – 3y = –11 2x – 6y = –14

OpenStudy (anonymous):

5x – 3y = –11 2x – 6y = –14 -- multiply this equation by 2.5 5x – 3y = –11 5x – 15y = –35-- subtract both equations 12y=24 y=2